OpenStudy (anonymous):

the bisector of angle B of an isosceles triangle ABC with AB=AC meets the circumcircle of triangle ABC at P if AP AND BC produced meet at Q prove that CQ=CA

OpenStudy (phi):



OpenStudy (phi):

triangle ABC is isosceles, with angle B = angle C Call the bisected angle at B = x (so angle B and angle C are both 2x) use these theorems: angle at Q formed by 2 secants = (angle of outer arc - angle of inner arc)/2 in other words angle Q= ( arc AB - arc PC)/2 also, inscribed angle PBC = (arc PC)/2. that is, x= arc PC /2 similarly 2x = arc AB/2

OpenStudy (phi):

you can now find angle Q in terms of x the angle QCA in triangle AQC is 180-2x you know Q, you know 180-2x so you can find the 3rd angle, angle CAQ you will find angle CAQ = angle AQC, so the triangle AQC is isosceles, with sides CQ= CA can you fill in the details?

OpenStudy (anonymous):

thank u so much @phi

OpenStudy (phi):

Here are the theorems, in case you have not seen them http://www.regentsprep.org/Regents/math/geometry/GP15/CircleAngles.htm the secant one used here is at the bottom of the page.

OpenStudy (anonymous):

k

OpenStudy (anonymous):



OpenStudy (anonymous):

also i don't understand this part, you can now find angle Q in terms of x the angle QCA in triangle AQC is 180-2x you know Q, you know 180-2x so you can find the 3rd angle, angle CAQ you will find angle CAQ = angle AQC, so the triangle AQC is isosceles, with sides CQ= CA can you fill in the details?

OpenStudy (phi):

Given isosceles triangle ABC with AB=AC that means angle ABC = angle ACB. I labeled the measure of these angles = 2x (that makes the measure of the bisected angle PBC = x )

OpenStudy (phi):

angle ACB is an "inscribed angle". It is equal to 1/2 the arc AB ACB= 2x, so we have: arc AB *1/2 = 2x and arc AB= 4x arc AP= 2x by the same reasoning angle Q= ( arc AB - arc PC)/2 angle Q= ( 4x - 2x)/2 = 2x/2 = x the angles in triangle AQC are : 180-2x, x, and a 3rd angle Y Y+x+180-2x = 180 Y= 180-180+x = x we see angle Y (that is, angle CAQ) has measure x. also, angle Q (that is, angle AQC) has measure x that means that triangle AQC is isosceles, with legs CA= CQ